User name Holmstrom
Log entry time 16:02:33 on July22,2003
Entry number 106237
keyword=Day Shift Summary
8:00 Continuing Kinematic 5.13
*** BHWP in, target 270deg
- run 2693/22693 07:53-08:10 3M CODA events
- run 2694/22694 08:11-08:26 3M CODA events
08:27 Called HallB + MCC to ask for the BHWP out
***BHWP out, target 270deg
- run 2695/22695 08:32-08:45 3M CODA events
- run 2696/22696 08:47-09:00 3M CODA events
08:58 MCC called, we will lose beam soon.
09:01 Beam down.
09:06 Changed target configuration from 270 deg. to 0 deg.
09:07 Ramped the Septum up to 280A.
09:11 Move Target to pickup Coil Position.
10:18 EPR-AFP loss test.
10:45 NMR-AFP loss test. Relative loss 0.53%/sweep.
11:00 Move Target to empty target position and set Septum to 450 A.
12:14 Set Septum Current at 390 A.
- run 22697
- run 2697/22698 11:33-12:48
- run 2699/22699 12:50-13:02
- run 2700/22700 13:13-13:56
- run 2701/22701 14:51-
12:20 Ramp Septum Current down to 300 A.
12:29 Ramp Septum Current down to 200 A.
12:40 Ramp Septum Current down to 148.5 A.
14:45 Request that MCC recover the following IOCs: iocha3 and iocha14.
15:03 Request that MCC have a software expert look at iocha3 and iocha14.
15:30 Ed did a hard reboot for the iocs including the left HRS magnets.
15:33 iocha14 is back up. iocha3 is not up, but it is the ioc for the RICH, which is not needed.
15:34 Cycling Q2 and Q3. Setting to 1600 A.
15:40 I did a hard reboot for the left HRS magnets ioc. It returned for about 3 minutes agian before crashing.
15:55 Javier Gomez called and said he would call Pam.